OpenStudy (shishi990):

solve the system using subtraction WILL FAN AND METAL 2x+2y=38 y=x+3

OpenStudy (bemoseemo):

Divide thru 1st by 2 x + y = 19 --- Substitute for "y" and solve for "x": x + x+3 = 19 2x = 16 x = 8 solve for "y": x + y = 19 8 + y = 19 y = 11

OpenStudy (bemoseemo):

your welcome.

OpenStudy (shishi990):

thank you

OpenStudy (bemoseemo):

I love algebra questions like this ¯\_(ツ)_/¯

OpenStudy (shishi990):

i dont get them... they are kinda hard for me

OpenStudy (bemoseemo):

if you need i can help you with a few more I have to go soon but at least 20 minutes

OpenStudy (shishi990):

ok thx

OpenStudy (bemoseemo):

np

OpenStudy (shishi990):

y=8-x 7=2-y

OpenStudy (shishi990):

@bemoseemo

OpenStudy (bemoseemo):

i answered that one on another site that is the link to it

OpenStudy (bemoseemo):

basically x= 13 y=-5

OpenStudy (shishi990):

ok thank you...!!!

OpenStudy (shishi990):

i passed thank you very much!!!